OpenStudy (anonymous):

| x-3|+| x-4| given that x>4 rewrite this expression without using absolute value notation

OpenStudy (across):

When x > 4, |x - 3| > 1 and x - 3 > 1. When x > 4, |x - 4| > 0 and x - 4 > 0. Therefore, when x > 4, |x - 3| + |x - 4| = x - 3 + x - 4 = 2x - 7 = 2(x - 3) - 1 > 1.

OpenStudy (across):

In other words, 2x - 7.